PHOTOASSOCIATIVE SPECTROSCOPY OF Cs_{2} MOLECULE

\underline{P. PILLET}, A. FIORETTI, D. COMPARAT, C. DRAG, A. CRUBELLIER, O. DULIEU, F. MASNOU-SEEUWS, and C. AMIOT, Laboratoire Aimé Cotton, CNRS II, Bat. 505, Campus d’Orsay, 91405 Orsay cedex, France.

We report the spectroscopy of molecular 1_{g}, 0_{u}^{+}, 0_{g}^{−} and 1_{u} state potentials below the dissociattion limit 6s^{2}S_{1/2} + 6p^{2}P_{3/2} by using the molecular photoassociation technique of cold Cs atoms, in a vapor-cell magneto-optical trap. The 1_{g} and 0_{u}^{+} states are detected by trap losses. The 0_{g}^{−} and 1_{u} states, which lead to the formation of cold molecules [1], are detected by photoionizing them into Cs_{2}^{+} and detecting with a time-of-flight mass spectrometer, while the 1_{g} and 0_{u}^{+} states are detected by trap losses. We give the spectra of the whole series 0_{g}^{−} and 1_{u}. The 0_{g}^{−} series is analyzed in a RKR approach and the hyperfine structure of the 1_{u} series will be discussed. The observed modulation of intensities of the lines is due to Franck Condon factors characterizing the photoassociative transition between the initial state of two free atoms and the finalro-vibrational level. The intensity modulation can give access to the scattering lengths of the ground state. This possibility will be discussed.

[1] A. Fioretti, D. Comparat, A. Crubellier, O. Dulieu, F. Masnou-Seeuws, P. Pillet, submitted to Phys. Rev. Lett.